OpenStudy (anonymous):

I was looking at a math team problem from 1998. The answer is 6.391. I'm interested in how to solve it. Thanks for the help: For the past three years, the inflation rate has been 4%, 3.17%, and 1.75%. For the past three years, income has increased at rates of 1.5%, 1.2% and 1.1%, respectively. If the inflation rate for this year is 1.2%, what income rate for this year would be necessary to match the rate of inflation for the past 4 years? Express answer to nearest 1000th of a percent.

OpenStudy (anonymous):

needs serious thinking.can u help me in telling from where u got this problem??

OpenStudy (anonymous):

6.32

OpenStudy (anonymous):

I'm a coach for a high school math team in Maine and this is an old question from a previous year that one of my students asked about.

OpenStudy (anonymous):

I just solved for \[\frac{4 + 3.17 + 1.75 + 1.2}{1.5+1.2+1.1+x} = 1\]

OpenStudy (anonymous):

which is why I was never allowed to join math teams at my high school :-P

OpenStudy (anonymous):

That's what I first thought to do but then when I checked the answer it should be 6.391.

OpenStudy (anonymous):

\[(1.04)\times (1.0317)\times (1.075)\times (1.012)=(1.015)\times (1.012)\times (1.011)\times (1+x)\] solve for x is my first guess

OpenStudy (anonymous):

that gives us 1.24030345997

OpenStudy (anonymous):

then i guess my first guess is wrong. now i have to think

OpenStudy (anonymous):

I will keep this question too, in order to annoy the math faculty at the uni :-P

OpenStudy (anonymous):

ok if i read the numbers correctly, an item that cost $1 4 years ago is now $1.10484587928 i got that from \[1.04 \times1.0317\times 1.0175 \times1.012\] are these numbers right?

OpenStudy (anonymous):

i get the same answer so i think this method is right

OpenStudy (anonymous):

solve \[(1.04)(1.0317)(1.0175)(1.012)=(1.015)(1.012)(1.011)(1+x)\] do get \[x=.0693078\] http://www.wolframalpha.com/input/?i=%281.04%29%281.0317%29%281.0175%29%281.012%29%3D%281.015%29%281.012%29%281.011%29%281%2Bx%29